OpenStudy (anonymous):

Find all values of x for which g'(x) does not exist. G(x)= |x+1|/(2x-4)

OpenStudy (anonymous):

certainly 2 is out, since it is not in the domain. also -1

OpenStudy (anonymous):

absolute values is a piecewise function, so you always have a problem where it is 0

OpenStudy (anonymous):

\[\frac{|x+1|}{2x-4} = \left\{ \begin{array}{lr} \frac{-x-1}{2x-4} & : x <0\\ \frac{x+1}{2x-4} & : x \geq 0 \end{array} \right.\]

OpenStudy (anonymous):

damn that is wrong

OpenStudy (anonymous):

and i spent all that time typing it. should say x < -1 and x > -1 i am not typing it over